answer all of the questions and show the solution  yrene Products manufactures recreational equipment. One of the companys products, a skateboard Over the past year the company sold 48,000 skateboards, with the following operating results Sales (48,000 skateboards) sells for $35 The skateboards are manufactured in an antiquated plant that relies heavily on direct labor workers. Thus, variable costs are high, totaling $24 50 per skateboard of which 70% direct labor cost S 1.680,000 1,176,000 Variable expenses Contribution margin 504,000 20,000 Fixed expenses Net operating income 84,000 Management is anxious to maintain and perhaps even improve its present level of income from the skateboards Required 1a. Compute the CM ratio and the break-even point in skateboards 1b. Compute the degree of operating leverage at last years level of sales 2. Due to an increase in labor rates, the company estimates that variable costs will increase by $1.75 per skateboard next year. If this change takes place and the selling price per skateboard remains constant at $35.00, what will be the new CM ratio and the new break-even point in skateboards? 3. Refer to the data in (2) above If the expected change in variable costs takes place, how many skateboards will have to be sold next year to earn the same net operating income, $84,000, as last year? 4. Refer again to the data in (2) above. The president has decided that the company may have to raise the selling price of its skateboards. If Tyrene Products wants to maintain the same CM ratio as last year, what selling price per skateboard must it charge next year to cover the increased labor costs? 5. Refer to the original data. The company is considering the construction of a new, automated plant The new plant would slash variable costs by 20%, but it would cause fixed costs to increase by 70% If the new plant is bult, what would be the companys new CM ratio and new break-even point in skateboards? 6. Refer to the data in (5) above a. If the new plant is built, how many skateboards will have to be sold next year to earn the same net operating income, $84,000, as last year? b-1.Assume that the new plant is constructed and that next year the company manufactures and sells 8,000 skateboards (the same number as sold last year). statement Prepare a contribution format income
Answer #1

1a) CM ratio= 504,000/1,680,000 =30%

Unit Sales to Breakeven= Fixed Cost/CM Per Unit

CM Per unit =504,000/48,000 =10.5

Unit Sales to Breakeven= 420,000/10.5

=40,000 Units

1b)Degree of Operating Leverage = 504,000/84,000 = 6

2)New CM Ratio= (35-26.25)/35

=25%

Unit Sales to Breakeven= 420,000/8.75

48,000 Units

3)Sales unit to Earn operating income 84,000 = (420,000+84,000)/8.75

=57,600 Units

4)New Sales Price to maintain CM Ratio of 30% =26.25/(1-30%)

=37.5

5)New variable cost per unit = 24.5 x(1-20%) =19.60

New fixed cost =420,000 x (1+70%)= 714,000

New CM Per unit =35-19.60= 15.40

CM Ratio= 15.40/35 =44%

Unit Sales to Breakeven =714,000/15.40

=46,363.6364 units

6a)714,000+84,00/15.40

=51818.1818

6b)

Contribution income Statement
Sales $ 1,680,000.00
variable Expenses $      940,800.00
Contribution Margin $      739,200.00
Fixed Expenses $      714,000.00
Net Operating Income $        25,200.00